OpenStudy (anonymous):

The area of a trapeazoid is A=0.5(b1+b2)h Find b1 if A=100, h=5, and b2=12

OpenStudy (anonymous):

\[The area of a trapeazoid is A=0.5(_{b1}+_{b2})h Find _{b1} if A=100, h=5, {b2}=12\]

OpenStudy (radar):

Replace the symbols with what you have. 100=.5 x 5(b1 + 12) 100= 2.5(b1+12) 100= 2.5b1 +30 2.5b1=100-30 2.5b1=70 b1=70/2.5 b1=28

OpenStudy (radar):

Did you follow that with understanding?

OpenStudy (radar):

to check: b1+b2=28+12=40 5 times 40 =200 1/2 time 200=100 A=100 Must be right,,,,,,agree or disagree.